# Bulk photovoltaic effect in ferroelectric and antiferroelectric phases of antimony sulphoiodide investigated by means of ab-initio simulations

### v1

@@ 4 Conclusions
In this paper, we presented a comprehensive investigation of the structural, electronic, and photoconductive properties of SbSI from first-principles, with particular attention to both the linear and circular photogalvanic contributions to the BPVE. Unlike previous studies, we centered our analysis on a comparison between the FE and the AFE phases.
We reproduced the crystal structures of the AFE and FE phases by using a hybrid functional approach combined with a van der Waals dispersion energy-correction term, which accounts for the weak interaction among the quasi-1D chains propagating along the $c$ axis. We found that, although the FE phase had a lower energy, the AFE phase was close in energy and therefore competitive. In the FE phase, the atoms of the two chains of the unit cell move in the same direction along the $c$-axis and a net polarization appears due to the shift of the cation sub-lattice with respect to the anion one. In contrast, in the AFE phase the cations of different chains move in opposite directions along $c$, leading to a different space group and to a zero net polarization.
@@
To explore the BPVE, the shift and the circular injection photoconductivity tensors were calculated in both phases, yielding different non-zero independent components in each phase, consistent with their distinct space groups. Physically, this implies that the direction of the photocurrent can change during the FE-AFE transition, providing a clear signature of the phase. Hence, we propose that the BPVE can serve as a reliable probe of the phase. Furthermore, we also showed that measuring the circular injection currents could access the spin-texture of the material, tracking its evolution from a Rashba-Dresselhaus to a Dresselhaus-Weyl character, and therefore, revealing the intricate coupling between structural and spintronic properties.
Finally, we predict that the shift photoconductivity of SbSI in the visible energy region is significantly larger than in most prototypical ferroelectric compounds, for both the FE and AFE phases. This finding confirms the potential of the material for photoconversion and photovoltaic applications.
@@ Acknowledgements
The authors acknowledge Julen Ibañez Azpiroz and Alvaro Ruiz Puente for their technical support in the calculation of the photoconductivities. This work was supported by the Next-Generation-EU program via the PRIN-2022 SORBET 1133 (Grant No. 2022ZY8HJY), the ICSC initiative (National Center for High-Performance Supercomputing, 1134 Big Data and Quantum Computing), the PE4 Partenariato Quantum Science and Technology (NQSTI). High-performance computing resources and support were provided by the CINECA under the ISCRA IsB28 “HEXTIM”, IsCc2 ”SFERA” and IsCa7 ”MUSCHI2D” projects.
@@
S6.3 Symmetry analysis for the photoconductivity tensors
@@ 3 $\sigma^{31}$ $\sigma^{32}$ $\sigma^{33}$ 0 0 0
S6.3.2 Circular injection photoconductivity

### v2

@@ 4 Conclusions
In this paper, we presented a comprehensive first-principles investigation of the structural, electronic, and photoconductive properties of SbSI, focusing in particular on both the linear and circular photogalvanic contributions to the BPVE. Our analysis compares in detail the FE and the AFE phases, as the phase transition occurs near room temperature, making both phases experimentally accessible.
We reproduced the crystal structures of the AFE and FE phases, elucidating the mechanisms responsible for their stabilization. In the FE phase, the atoms of the two chains of the unit cell shift in the same direction along the $c$-axis, resulting in a net polarization due to the shift of the cation sub-lattice with respect to the anion one. In contrast, in the AFE phase, the cations in adjacent chains move in opposite directions along $c$, leading to a zero net polarization.
@@
To explore the BPVE, we performed a comprehensive analysis of both the shift and circular injection photoconductivity tensors, identifying all symmetry-allowed components and accurately estimating their magnitudes. The results can serve as a practical guide for distinguishing between the different phases of SbSI through photocurrent measurements. Moreover, we predict that the shift photoconductivity of SbSI in both phases can reach peak values significantly larger than those of most prototypical FE compounds, far exceeding those reported in experiments to date. This suggests the potential to improve the material’s photovoltaic performance beyond the current state of the art.
Finally, we demonstrated that circular injection currents can serve as sensitive probes of the changes in spin texture, revealing a strong coupling between structural, optical, and spintronic properties. This finding underscores the promise of SbSI as a platform for exploring fundamental aspects of multifunctional behavior.
@@ Acknowledgements
The authors acknowledge Julen Ibañez Azpiroz and Alvaro Ruiz Puente for their technical support in the calculation of the photoconductivities. This work was supported by the Next-Generation-EU program via the PRIN-2022 SORBET 1133 (Grant No. 2022ZY8HJY), the ICSC initiative (National Center for High-Performance Supercomputing, 1134 Big Data and Quantum Computing), the PE4 Partenariato Quantum Science and Technology (NQSTI). High-performance computing resources and support were provided by the CINECA under the ISCRA IsB28 “HEXTIM”, IsCc2 "SFERA" and IsCa7 "MUSCHI2D" projects.
@@
Appendix S1 Symmetry mode analysis for the AFE structure
Figure S1: Energy as a function of AFE $\Gamma_{1}^{-}$ distortion, calculated by using the GGA (red line) and HSE (black line).
The symmetry mode analysis reveals that the $\Gamma_{1}^{-}$ mode is the primary distortion driving the transition from the PE to the AFE phase in SbSI. Starting from the centrosymmetric PE Pnam structure, we introduce various amplitudes of the $\Gamma_{1}^{-}$ distortion, keeping the lattice parameters fixed, and compute the total energies of the resulting structures.
Our HSE calculations show a clear energy gain relative to the PE phase as the $\Gamma_{1}^{-}$ distortion is condensed, thus providing a compelling energetic justification for the structural transition to the AFE phase. In contrast, analogous calculations performed using the standard GGA functional indicate an energy increase, which fails to stabilize the AFE structure.
@@
S7.3 Symmetry analysis for the photoconductivity tensors
@@ 3 $\sigma^{31}$ $\sigma^{32}$ $\sigma^{33}$ 0 0 0
S7.3.2 Circular injection photoconductivity